Alternate integrals signs with L^{A}TEX 2ε

Eddie Saudrais

version 1.1 01/20/2005

Abstract

The package esint.sty allows you to use new integrals symbols.

1 Installation

Run L^{A}TEX 2ε on esint.ins to generate files:

1. Put esint.sty on TEXINPUT.

2. Put uesint.fd on TEXINPUT, for example with esint.sty.

3. Put esint10.mf, mathint.mf and bigint.mf on MFINPUT.

Run on esint10.mf file to generate esint10.tfm files:

mf \mode=localfont; input esint10.mf

Put esint10.tfm on the right place.

2 Using esint

Load the package with \usepackage{esint}, and enjoy! BE CAREFUL: if you are using amslatex, you must load esint AFTER

amslatex.

Available integrals signs:

Commande textstyle displaystyle

    \int  ∫  ∫

   \iint  ∬  ∬

   \iiint  ∭  ∭

  \iiiint  ⨌  ⨌

  \dotsint  ∫⋯∫  ∫⋯∫

 \oint  ∮ ∮ \oiint  ∯ ∯ \varoiint  ∯ ∯ \sqint  ⨖ ⨖ \sqiint  ⨖⨖ ⨖⨖ \ointctrclockwise  ∳ ∳ \ointclockwise  ∲ ∲ \varointclockwise  ∲ ∲ \varointctrclockwise ∳ ∳ \fint  ⨏ ⨏ \landupint  ⨙ ⨙ \landdownint  ⨚ ⨚ You can customize the space between in-

tegral sign in multiple integrals. You have to modify lines 12 and 13 of the esint10.mf file: tdec# and ddec# are spaces between signs. If you modify esint10.mf, delete esint10.tfm, the generated *.pk files, and run

on esint10.mf.

3 Updates

• 20/01/2005: change in esint.fd in order to avoid a problem inside align environment. Thank’s to Eckhard Neber. Font files (mf, pfb, tfm...) are unchanged.

4 The code

The package identifies himself

1 ⟨∗package⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{esint}

To redifine symbols

 4 \def\re@DeclareMathSymbol#1#2#3#4{%
 5  \let#1=\undefined
 6  \DeclareMathSymbol{#1}{#2}{#3}{#4}}

Definition of the symbol font:

 7 \DeclareSymbolFont{largesymbolsA}{U}{esint}{m}{n}

Definition of the new symbols:

 8 \re@DeclareMathSymbol{\intop}{\mathop}{largesymbolsA}{'001}
 9  \def\int{\intop\nolimits}
10 \re@DeclareMathSymbol{\iintop}{\mathop}{largesymbolsA}{'003}
11  \def\iint{\iintop\nolimits}
12 \re@DeclareMathSymbol{\iiintop}{\mathop}{largesymbolsA}{'005}
13  \def\iiint{\iiintop\nolimits}
14 \re@DeclareMathSymbol{\iiiintop}{\mathop}{largesymbolsA}{'007}
15  \def\iiiint{\iiiintop\nolimits}
16 \re@DeclareMathSymbol{\dotsintop}{\mathop}{largesymbolsA}{'011}
17  \def\dotsint{\dotsintop\nolimits}
18 \re@DeclareMathSymbol{\ointop}{\mathop}{largesymbolsA}{'013}
19  \def\oint{\ointop\nolimits}
20 \re@DeclareMathSymbol{\oiintop}{\mathop}{largesymbolsA}{'015}
21  \def\oiint{\oiintop\nolimits}
22 \re@DeclareMathSymbol{\sqintop}{\mathop}{largesymbolsA}{'017}
23  \def\sqint{\sqintop\nolimits}
24 \re@DeclareMathSymbol{\sqiintop}{\mathop}{largesymbolsA}{'021}
25  \def\sqiint{\sqiintop\nolimits}
26 \re@DeclareMathSymbol{\ointctrclockwiseop}{\mathop}{largesymbolsA}{'027}
27  \def\ointctrclockwise{\ointctrclockwiseop\nolimits}
28 \re@DeclareMathSymbol{\ointclockwiseop}{\mathop}{largesymbolsA}{'031}
29  \def\ointclockwise{\ointclockwiseop\nolimits}
30 \re@DeclareMathSymbol{\varointclockwiseop}{\mathop}{largesymbolsA}{'033}
31  \def\varointclockwise{\varointclockwiseop\nolimits}
32 \re@DeclareMathSymbol{\varointctrclockwiseop}{\mathop}{largesymbolsA}{'035}
33  \def\varointctrclockwise{\varointctrclockwiseop\nolimits}
34 \re@DeclareMathSymbol{\fintop}{\mathop}{largesymbolsA}{'037}
35  \def\fint{\fintop\nolimits}
36 \re@DeclareMathSymbol{\varoiintop}{\mathop}{largesymbolsA}{'041}
37  \def\varoiint{\varoiintop\nolimits}
38 \re@DeclareMathSymbol{\landupintop}{\mathop}{largesymbolsA}{'043}
39  \def\landupint{\landupintop\nolimits}
40 \re@DeclareMathSymbol{\landdownintop}{\mathop}{largesymbolsA}{'045}
41  \def\landdownint{\landdownintop\nolimits}

42 ⟨/package⟩

43 ⟨∗fdfile⟩

Font definition file:

44 \ProvidesFile{uesint.fd}
45 \DeclareFontFamily{U}{esint}{}
46 \DeclareFontShape{U}{esint}{m}{n}{
47 <-> esint10
48 }{}

49 ⟨/fdfile⟩

4